# MVCNet: Multi-View Contrastive Network for Motor Imagery Classification

### v3

@@ IV-E Comparison with Data Augmentation Strategies
  3. 3.
    MVCNet consistently outperforms all augmentation strategies across all CNN backbones. This indicates that MVCNet integrates more informative transformations and learns more robust and discriminative EEG representations under the challenging cross-subject scenario.
@@ IV-F Effect of Multi-View Augmentation
To investigate the impact of combining different augmentation views, we evaluated all pairwise combinations among the seven augmentation strategies using EEGNet as the backbone. As shown in Fig. 6, combinations of two views generally yield better performance than individual augmentations, demonstrating the complementary benefits of multi-view augmentation.
@@ IV-G Feature Visualization
$t$-SNE [49] visualizations were conducted on features extracted from EEGNet and MVCNet under different augmentation strategies on the Blankertz2007 dataset, as shown in Fig. 7. Compared with EEGNet, the feature clusters in MVCNet are more compact, with samples from different augmentation views better aligned within each class. This demonstrates the effectiveness of our cross-view contrastive module in bridging the gaps among time, spatial, and frequency domain transformations.
@@ IV-H Ablation Study and Parameter Sensitivity Analysis
To further assess the robustness of the proposed approach, sensitivity analyses were performed on the weighting parameters $\lambda$ and $\gamma$. In Fig. 8, one parameter was varied while the other was fixed at 0.1. Observe that the performance remained stable across a wide range of values.
@@ V Conclusion
This paper proposed MVCNet, a dual-branch network for MI decoding. By integrating CNN and Transformer modules in parallel, MVCNet effectively captures both local spatial-temporal features and global temporal dependencies. Cross-view and cross-model contrastive modules were introduced to enforce feature consistency across multiple views and network branches. Experimental results on five public MI datasets under three evaluation scenarios demonstrated that MVCNet consistently outperformed existing seven state-of-the-art models. Future research will explore more informative and effective model architectures, such as Mamba-based hybrid models, to further improve MI decoding performance.

### v4

@@ IV-E Comparison with Data Augmentation Strategies
  • MVCNet outperforms most augmentation strategies across all CNN backbones, obtaining the best or second-best performance. This indicates that MVCNet integrates more informative transformations and learns more robust and discriminative EEG representations under the challenging cross-subject scenario.
@@ IV-F Effect of Multi-View Augmentation
To investigate the impact of combining different augmentation views, we evaluated all pairwise combinations among the seven augmentation strategies using EEGNet as the backbone. As shown in Figure 6, combinations of two views generally yield better performance than individual augmentations, demonstrating the complementary benefits of multi-view augmentation.
@@ IV-G Feature Visualization
$t$-SNE [54] visualizations were conducted on features extracted from EEGNet and MVCNet under different augmentation strategies on the Blankertz2007 dataset, as shown in Figure 7. Compared with EEGNet, the feature clusters in MVCNet are more compact, with samples from different augmentation views better aligned within each class. This demonstrates the effectiveness of our cross-view contrastive module in bridging the gaps among time, spatial, and frequency domain transformations.
@@ IV-H Ablation Study and Parameter Sensitivity Analysis
To further assess the robustness of the proposed approach, sensitivity analyses were performed on the weighting parameters $\lambda$ and $\gamma$. In Figure 8, one parameter was varied while the other was fixed at 0.1. Observe that the performance remained stable across a wide range of values.
@@
IV-I Effect of Transformer Heads and Layers
To evaluate the effect of the number of Transformer heads and layers, we conducted an additional ablation study by varying each parameter while keeping the other fixed. Specifically, we varied the number of Transformer layers $N_{L}\in\{1,2,3,4\}$ with a fixed number of heads $N_{H}=2$, and varied $N_{H}\in\{1,2,4,8\}$ with $N_{L}=2$ on BNCI2014001 dataset. Results are summarized in Figure 9. Observed that:
  • For the number of layers, increasing $N_{L}$ from 1 to 2 improves accuracy, but further increases to 3 or 4 do not yield additional gains and may even lead to slight degradation. This suggests that deeper attention structures may introduce overfitting or optimization challenges in small-scale EEG datasets.
  • For the number of heads, accuracy improves as $N_{H}$ increases from 1 to 2 or 4, indicating that multi-head attention enhances the model’s ability to capture diverse temporal patterns. However, setting $N_{H}=8$ leads to a slight performance drop, possibly due to redundant subspace partitioning or insufficient training data to support larger attention capacity.
  • Overall, the better performance is achieved with $N_{L}=2$ and $N_{H}=2$, which we adopt as the default configuration for MVCNet throughout the main experiments.
Figure 9: Ablation study on the number of Transformer layers $N_{L}$ and heads $N_{H}$.
@@ V Conclusion
This paper proposed MVCNet, a dual-branch network for MI decoding. By integrating CNN and Transformer modules in parallel, MVCNet effectively captures both local spatial-temporal features and global temporal dependencies. Cross-view and cross-model contrastive modules were introduced to enforce feature consistency across multiple views and network branches. Experimental results on five public MI datasets under three evaluation scenarios demonstrated that MVCNet consistently outperformed existing nine state-of-the-art models. Future research will explore more informative and effective model architectures, such as Mamba-based hybrid models, to further improve MI decoding performance.